# Quantum metrology through spectral measurements in quantum optics

### v1

@@ H.2 Extension to $N$ bosonic modes
This procedure can be generalized to any number of modes. For an $N$-mode system $\mathcal{H}=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$], the joint distribution $P_{n_{1},n_{2},\ldots,n_{N}}$ can be reconstructed via a multi-dimensional extension of the same formula:
@@
Here, in this section, we are going to prove the inequality presented in Eq. (18), establishing that $F_{\theta}[p(n_{1},n_{2}|\theta)]\geq 1/2(F_{\theta}[p(n_{1}|\theta)]+F_{\theta}[p(n_{2}|\theta)])$. To prove this, we consider the case of a general joint density matrix $\hat{\rho}_{\theta}$ describing the quantum state of both bosonic sensors. Although in practice the Hilbert space of these sensors is typically restricted to a few excitations, we will consider here the general case up to $(N_{1},N_{2})$-excitations. For the shake of simplicity, we also consider the particular case of standard photon-counting $\vec{\alpha}=\vec{0}$. The computation of the Fisher information in Eq. (5) requires the use of the joint probability distribution, $p(n_{1},n_{2}|\theta)$, with matrix representation

### v2

@@ H.2 Extension to $N$ bosonic modes
This procedure can be generalized to any number of modes. For an $N$-mode system ($\mathcal{H}=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$), the joint distribution $P_{n_{1},n_{2},\ldots,n_{N}}$ can be reconstructed via a multi-dimensional extension of the same formula:
@@
Here, in this section, we are going to prove the inequality presented in Eq. (22), establishing that $F_{\theta}[p(n_{1},n_{2}|\theta)]\geq 1/2(F_{\theta}[p(n_{1}|\theta)]+F_{\theta}[p(n_{2}|\theta)])$. To prove this, we consider the case of a general joint density matrix $\hat{\rho}_{\theta}$ describing the quantum state of both bosonic sensors. Although in practice the Hilbert space of these sensors is typically restricted to a few excitations, we will consider here the general case up to $(N_{1},N_{2})$-excitations. For the sake of simplicity, we also consider the particular case of standard photon-counting $\vec{\alpha}=\vec{0}$. The computation of the Fisher information in Eq. (8) requires the use of the joint probability distribution, $p(n_{1},n_{2}|\theta)$, with matrix representation